Write spin multiplicity of spin S=1/2 particle.
1
Expert's answer
2020-11-05T07:33:10-0500
"s_1=\frac{1}{2},s_2=-\frac{1}{2}\to s=0,2s+1=1\\\text{ (singlet state)}\\s_1=\frac{1}{2},s_2=\frac{1}{2}\to s=1,2s+1=3\\\text{ (triplet state)}"